If a and b are prime numbers which if the following cannot be the sum of a and b?
A)13
B)16
C)25
D)27

Answer by ikleyn(53765)

.
13 = 2 + 11.


16 = 5 + 11.


25 = 2 + 23.


27 can not be the sum of two primes, since then one of them must be even number (i.e. 2),  but 27 - 2 = 25 IS NOT a PRIME.


ANSWER.  27.